Program

ECCAD’2001

May 5, Tallahassee, FL

8:00 – 8:45 Registration. 845 – 9:00 Opening Remarks. 9:00 – 10:00 Invited Lecture: Arne Storjohann (U. of Western Ontario).

Design and Analysis of Algorithms for Symbolic Linear Algebra.

10:00 – 10:30 Break. 10:30 – 11:30 Invited Lecture: Hoon Hong (NCSU).

Computer Algebra for Composed Polynomials.

11:30 – 1:30 Lunch. 1:30 – 2:30 Poster Session 1. 2:30 – 3:30 Invited Lecture: Alexander Hulpke (Ohio State U.).

Working with Quotients of Finitely Presented Groups.

3:30 – 4:30 Poster Session 2. 4:30 Closing Remarks. 5:00 – 6:00 Reception at the Turnbull Center. 6:00 – 8:00 Dinner Buffet will be served at the Turnbull Center.

Poster sessions ECCAD’2001 May 5, Tallahassee, FL

Poster Session 1.

• Mhenni Benghorbal Approximate Solution Of The Fractional Differential Equation D^{q}y(x) = λy(x) (joint work with Robert Corless)

• Nikolaj Glazunov Computer Algebra of Zeta Functions and Distribution of Numerical Sequences

• Markus Hitz Penrose Revisited

• Claude-Pierre Jeannerod Computing matrix perturbations with minimal leading Jordan structure

• Jeremy Johnson SPIRAL: A System for Implementation and Platform-Adaptation of Signal Processing Algorithms (joint work with Markus Pueschel)

• Ilias Kotsireas Implicitization of curves surfaces and hypersurfaces

• Werner Krandick Parallel Isolation of Polynomial Real Roots.

Poster Session 2.

• John May Gao’s Bivariate Polynomial Factorisation Algorithm: A Maple Implementation

• Yuri Movsisyan Hyperidentities And Hypervarieties, classification, characterization and application

• Jack Perry Groebner Bases for Composed Polynomials (joint work with Hoon Hong).

• Andreas Strotmann Intelligent Mathematics Tutoring and Testing: Two Experiments and their Lessons

• Vladimir Tonchev A mass formula for Steiner triple systems ST S(2^{n} − 1) of 2-rank 2^{n} − n

• William Turner A Randomized Baby Steps/Giant Steps Implementation of Wiedemann’s Determinant Algorithm (joint work with Erich Kaltofen).

• Xiaofang Xie Symbolic Circuit Analysis in Maple (joint work with Robert Corless and Stephen Watt)

• Eugene Zima On computational properties of chains of recurrences.